# arXiv:2603.00222v1[cs.CR] 27 Feb 2026

### v1

@@ Abstract
The Forensics Investigations Network in Digital Sciences (FINDS) Research Center of Excellence (CoE) is a pioneering research and education initiative funded by the US Army Research Laboratory. FINDS conducts innovative research, fosters cutting-edge workforce development programs, and bridges the intersection of science education and technology. The workforce development efforts of the CoE have played a critical role in strengthening national defense by training skilled professionals and devising solutions to bolster the resilience of critical infrastructures and businesses against modern cyber-attacks. The emerging field of Digital Forensic Engineering Education is at the forefront of addressing the pressing challenges of equipping the next generation of cyber warriors.
This paper highlights the core components of FINDS, presenting a comprehensive framework for workforce development and training aimed at empowering underrepresented and non-traditional students in cybersecurity and digital forensic engineering. Building on a legacy of successful initiatives targeting underrepresented groups, the approach expands participation and inclusivity in cybersecurity engineering research and digital forensics technology.
Central to this paper are strategies for capacity building through the Multidependency Capacity Building Skills Graph data structure. Key focus areas include AI-powered digital forensic programming, high-performance computing, and curriculum development integrated with software development training. Decision Tree Classifier and Statistical Modeling, are employed to accurately determine parameters and leverage real-time strategy datasets encompassing peer mentoring, curriculum modules, hands-on community technologies, cybersecurity workshops, and entrepreneurial extension services. Rigorous statistical experiments conducted over three years on capacity evaluation datasets have provided robust insights. The analysis of data and student feedback underscores the outstanding results achieved in capacity building within Digital Forensic Engineering Education.

### v2

# Empowering Future Cybersecurity Leaders: Advancing Students through FINDS Education for Digital Forensic Excellence

@@ Abstract
The Forensics Investigations Network in Digital Sciences (FINDS) Research Center of Excellence (CoE), funded by the U.S. Army Research Laboratory, advances Digital Forensic Engineering Education (DFEE) through an integrated research–education framework for AI-enabled cybersecurity workforce development. FINDS combines high-performance computing (HPC), secure software engineering, adversarial analytics, and experiential learning to address emerging cyber and synthetic media threats.
This paper introduces the Multidependency Capacity Building Skills Graph (MCBSG), a directed acyclic graph–based model that encodes hierarchical and cross-domain dependencies among competencies in AI-driven forensic programming, statistical inference, digital evidence processing, and threat detection. The MCBSG enables structured modeling of skill acquisition pathways and quantitative capacity assessment.
Supervised machine learning methods, including entropy-based Decision Tree Classifiers and regression modeling, are applied to longitudinal multi-cohort datasets capturing mentoring interactions, laboratory performance metrics, curriculum artifacts, and workshop participation. Feature importance analysis and cross-validation identify key predictors of technical proficiency and research readiness.
Three-year statistical evaluation demonstrates significant gains in forensic programming accuracy, adversarial reasoning, and HPC-enabled investigative workflows. Results validate the MCBSG as a scalable, interpretable framework for data-driven, inclusive cybersecurity education aligned with national defense workforce priorities.